OpenStudy (anonymous):

if (cosx)^2 = (1+cos2x)/2 then (secx)^2=2/(1+cos2x) true?

OpenStudy (lalaly):

yeh

OpenStudy (anonymous):

this question is telling me to "complete the identity" of (secx)^2-(tanx)^2=? any thoughts?

OpenStudy (anonymous):

i can put them all in terms of cosine and fig it out, yeh?

OpenStudy (lalaly):

sec^2x=1+tan^2x

OpenStudy (anonymous):

damn you're good

OpenStudy (lalaly):

lol :D

OpenStudy (anonymous):

so 1+tan^2x-tan^2x =..... 1? LOL

OpenStudy (lalaly):

obviously... the tan^2x cancels nd ur left wit 1 so 1=1 thats right.... the answer is (secx)^2-(tanx)^2=1

OpenStudy (anonymous):

sorry. did the ... to imply i'm an idiot, pre-obviously :D

OpenStudy (lalaly):

u can derive all that from sin^2x+cos^2x=1 if u divide them all by cos^2x u get the equation i wrote